# Natural Language Dataset Generation Framework for Visualizations Powered by Large Language Models

### v1

@@ 6. Discussion
Figure 4. Four examples of generated L1/L2 captions with corresponding charts. We found that our framework can successfully generate captions even on complex charts with varying interactions and multiple views.
6.1. Strengths and Weaknesses of LLM Framework
6.1.1. The Framework Can Guide Itself via Key Questions
@@
6.1.2. The Framework Works Robustly on Different Chart Complexity
@@
6.1.3. The Framework Depends Highly on Vega-Lite Specifications
@@
6.1.4. The Framework Predicts Only Common Chart Types
@@ 6.2. Limitations and Future Work
6.2.1. Enriching Capabilities of the Framework through External Information or Human-in-the-Loop Approach
While Vega-Lite specifications serve as powerful inputs for generating various types of NL datasets, it is inherently challenging to extract information that does not exist within these specifications. Considering that our framework operates in a fully automatic manner, it does not rely on any external resources. However, this limitation can potentially impact the performance of NL generation, as it aligns with observations in guided discovery, where insufficient prior knowledge can hinder learners from formulating hypotheses, interpreting data effectively, and engaging in systematic experimentation (De Jong and Van Joolingen, 1998).
To enhance the capabilities of our LLM framework, we suggest accessing external sources of information or adopting a human-in-the-loop approach to guide the process during NL dataset generation. For instance, generating L3/L4 captions often necessitates access to common or domain-specific knowledge (Lundgard and Satyanarayan, 2021). In this regard, employing tools like ReAct (Yao et al., 2023) becomes advantageous, as it facilitates reasoning to assist the model in deducing, tracking, and updating action plans while also handling exceptions. This enables us to proactively retrieve information from the web when required. Similarly, in the context of providing guidance, domain experts can contribute by specifying particular directions they wish to proceed. By providing key questions, they can guide the model and steer its focus accordingly. This collaborative approach between automated NL generation and human expertise offers the potential to address knowledge gaps and improve the quality and relevance of generated NL datasets.
@@ 7. Conclusion
In this study, we introduce an LLM framework designed to generate diverse NL datasets aimed at enhancing NLIs for data visualization research. Our framework takes a Vega-Lite specification as only input and employs multiple prompting techniques to accurately generate various NL datasets, including captions, utterances, and questions. We also propose a score-based paraphrasing approach to enhance the syntactic diversity of the generated NL datasets. To demonstrate the robust and generalizable performance of our framework, we present a dataset comprising 1,981 Vega-Lite specifications. This dataset surpasses the baselines in terms of complexity and diversity. Our experimental results substantiate that the framework excels in accurately generating both L1 and L2 captions, while achieving higher diversity in the generation of utterances and questions compared to the baselines. We hope our framework and chart dataset can advance research in developing diverse NLIs for data visualization.
Appendix A Prompt for NL Generation Pipeline
@@
  • vl [ALL]: Pre-processed Vega-Lite specification converted as one line string;
  • ftt_str [L1, L2, Utterance]: Information about fields, titles, types, and values;
  • prompt [L2]: Questions derived from the guided discovery process;
  • info [L2]: Answers for the questions computed through LangChain library (Chase, 2022);
@@ A.2. L2 Caption
5Generate a concise level 2 natural language description of a visualization, with 1 or 2 sentences for each level.
6The level 2 description should contain statistical and relational properties of the visualization (e.g., descriptive statistics, extrema, outliers, correlations, point-wise comparisons).
7
8##
9Level 2 NL Description:
@@ A.4. Question
12Step 8. Generate a compositional question using this value, without including any visual attributes such as color, length, size, or position.
@@ Appendix B Prompt for Automatic Qualitative Coding
Appendix C Gold Reference Guidelines
C.1. Utterance

### v3

@@ 7. Discussion
Figure 5. Four examples of generated L1/L2 captions with corresponding charts. We found that VL2NL can successfully generate captions even on complex charts with varying interactions and multiple views.
7.1. Strengths and Weaknesses of VL2NL
7.1.1. VL2NL Can Guide Itself via Key Questions
@@
7.1.2. VL2NL Works Robustly on Different Chart Complexity
@@
7.1.3. VL2NL Depends Highly on Vega-Lite Specifications
@@
7.1.4. VL2NL Predicts Only Common Chart Types
@@ 7.2. Limitations and Future Work
7.2.1. Enriching Capabilities of VL2NL through External Resources
While Vega-Lite specifications serve as powerful inputs for generating various types of NL datasets, it is inherently challenging to extract information that does not exist within these specifications. Although our framework can operate in both fully-automatic and mixed-initiative manner, it does not rely on external resources. This limitation can potentially impact the performance of NL generation, as it aligns with observations in guided discovery, where insufficient prior knowledge can hinder learners from formulating hypotheses, interpreting data effectively, and engaging in systematic experimentation (De Jong and Van Joolingen, 1998). To enhance the capabilities of VL2NL, we suggest accessing external information to guide the process during NL dataset generation. For instance, generating L3/L4 captions often necessitates access to common or domain-specific knowledge (Lundgard and Satyanarayan, 2021). In this regard, employing tools like ReAct (Yao et al., 2023) becomes advantageous, as it facilitates reasoning to assist the model in deducing, tracking, and updating action plans while also handling exceptions. This enables us to proactively retrieve information from the web when required.
@@ 7.2.2. Augmenting Vega-Lite Specifications
7.2.3. Covering Additional NL datasets
Our framework exhibits potential for generalization across multiple NL datasets. However, we recognize that it covers only limited number of types, which we aim to expand in our future research. Specifically, we plan to create conversational NL datasets to facilitate interactive communication with NLIs, given the growing significance of conversational agents. We believe a dataset based on deeper analysis of users’ conversational characteristics will be immensely beneficial for researchers. We also plan to address reference datasets linking charts with text to help make interactive documents. We believe this will make the connection between them clearer, and the reading experience more enjoyable and engaging.
@@ 8. Conclusion
We introduce VL2NL designed to generate diverse NL datasets aimed at enhancing NLIs for data visualization research. Our framework takes a Vega-Lite specification as only input and employs guided discovery to accurately generate various NL datasets, including captions, utterances, and questions. We also propose a score-based paraphrasing approach to enhance the syntactic diversity of the generated NL datasets. We also present a dataset comprising 1,981 Vega-Lite specifications. This dataset surpasses the baselines in terms of complexity and diversity. Our experimental results substantiate that the framework excels in accurately generating both L1 and L2 captions, while achieving higher diversity in the generation of utterances and questions compared to the baselines. Last, we introduce real-world scenarios of using LLM-generated NL datasets and our framework. We hope our framework and chart collection can advance research in developing NLIs for data visualization.
Appendix A Prompts for NL Generation
@@
  • vl [all]: Minified Vega-Lite specification;
  • ftt_str [L1/L2 caption, utterance]: Information about fields, titles, types, and values;
  • prompt [L2 caption]: Questions derived from the guided discovery process;
  • info [L2 caption]: Answers for the questions computed through LangChain library (Chase, 2022);
@@ A.2. L2 Caption
5Generate a concise level 2 NL description of a visualization, with 1 or 2 sentences. It contains statistical and relational properties of the visualization (e.g., descriptive statistics, extrema, outliers, correlations, point-wise comparisons).
6
7##
8Level 2 NL Description:
@@ A.4. Question
12Step 8. Generate a compositional question using these operations, without including any visual attributes such as color, length, size, or position.
@@ Appendix B Prompt for Automatic Qualitative Coding
Appendix C Prompts for Score-based Paraphrasing
We explain the variables used in our prompts:
  • Example Sentence: A sentence we want to paraphrase;
  • Axis: An explanation about each of the defined language axes;
  • Direction: A set of two opposite directions of the given language axis;
  • Score: A specific value on a Likert-scale ranging from one to five assigned to each of the language axis.
C.1. Paraphrasing with one axis
1{Axis}
2
4
5Sentence: {Example Sentence}
6Score: {Score}
C.2. Paraphrasing with two axes
1{Axis-1}
2{Axis-2}
3
7
8Sentence: {Example Sentence}
9Score-A: {Score-A}, Score-B: {Score-B}
Appendix D Gold Reference Guidelines
D.1. Utterance